OpenStudy (anonymous):

What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. –5 m/s2 B. +5 m/s2 C. –25 m/s2 D. +25 m/s2

OpenStudy (anonymous):

@alphadxg

OpenStudy (alphadxg):

(final velocity - initial velocity) -------------------------------------- (final time - initial time) = ( 0 - 25 ) m/s / ( 5.0 - 0 ) s = -25/5 = -5 m/s^2 @cool41

OpenStudy (anonymous):

so it would be d? @alphadxg

OpenStudy (anonymous):

no mevermind a